For which value of k will the vector v1=(1, −2, k) in R



3



be a linear combination of



v2 = (3, 0, −2) and v2= (2, −1, 5)?

1
Expert's answer
2022-02-23T08:20:50-0500
"\vec v_1=a\vec v_2+b\vec v _3"

"3a+2b=1"

"0-b=-2"

"-2a+5b=k"

"a=-1, b=2, k=12"

"k=12"